You deposit $300 in a bank account that pays an interest rate of 6%, compounded quarterly. In excel, compute how much your account balance will be at the end of each year for the next 5 years. Which of the following accurately depicts this calculation?
Answer #1

D 2 A 1 PV 2 Rate 3 Rate/Quarter 4. Period 5 Period (quarte 6 FV 7 B C $300 6% annually 1.50% 5 years 20 quarters $404.06B 300 0.06 annually =B2/4 А 1 PV 2 Rate 3 Rate/Quarter 4 Period 5 Period (quarters) 6 FV 7 8 years quarters =B4*4 =FV(B3,B5,,